OpenStudy (anonymous):

The given lengths are two sides of a right triangle. All three side lengths of the triangle are integers and together form a pythagorean triple. Find the length of the third side, and tell whether it is a leg or a hypotenuse. 24, 51 A. 21 leg B. 45 leg C. 48 leg D. 56 hypotenuse

OpenStudy (ranga):

a^2 + b^2 = c^2 Try a = 24, b = 51 and find out if c will be an integer. If not, try a = 24, c = 51 and find out if b will be an integer.